The volume of a gas changes from 4.5 L to 3 L when we lower the temperature to 250 K. What must have been the initial temperature in K?
1
Expert's answer
2020-06-03T15:16:52-0400

Solution: As a consequence of the law of Mendeleev-Clapeyron, we have: V1/T1=V2/T2. By condition:

V2=3 L; T2=250 K; V1=4.5 L, then T1=V1*T2/V2=4.5*250/3=375 K.

Answer: 375 K.